# Joint Transmit and Jamming Power Optimization for Secrecy in Energy Harvesting Networks: A Reinforcement Learning Approach

### v1

@@ VII Results and Discussions
Fig. 3(a) plots the expected total discounted reward versus the EH probability while $p=q$ and compares the performance of algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency plot can be found in Fig. 3(b). From Fig. 3(a) we find that the expected total discounted reward increases for all of the algorithms as EH probability increases. A higher probability of EH leads to more energy available at the nodes which leads to more expected total transmitted secure bits until the network stops functioning, hence the observation. However, more expected total transmitted secure bits does not mean more energy efficiency as energy efficiency decreases with the increasing EH probability in 3(b). As energy efficiency is the ratio of total transmitted secure bits and total energy expenditure in the system, it seems for the algorithms OJPA, RSJPA, and GA, the rate of increase in the total transmitted secure bits and the rate of increase in the energy expenditure are not the same. This suggests that the optimization for total transmitted secure bits alone without taking energy efficiency into account is not beneficial. Rather one needs to optimize the system taking both into account. The observation and the corresponding reasoning for the NA is just the opposite.
@@
Figs. 4(a) and 4(c) plot expected total discounted reward versus $B_{\textrm{S}}^{\textrm{max}}$ and $B_{\textrm{D}}^{\textrm{max}}$, respectively and compares the performance of algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency plots are shown in Figs. 4(b) and 4(d), respectively. We observe that as $B_{\textrm{S}}^{\textrm{max}}$ or $B_{\textrm{D}}^{\textrm{max}}$ increases, the expected total discounted reward improves. When $B_{\textrm{S}}^{\textrm{max}}$ and $B_{\textrm{D}}^{\textrm{max}}$ is larger, nodes S and D can store more harvested energy which leads to more expected total transmitted secure bits. If we consider energy efficiency in Figs. 4(b) and 4(d), energy efficiency also improves when $B_{\textrm{S}}^{\textrm{max}}$ increases, however, the same observation is not true when $B_{\textrm{D}}^{\textrm{max}}$ increases. When $B_{\textrm{D}}^{\textrm{max}}$ increases, energy efficiency first improves, however, later degrades as $B_{\textrm{D}}^{\textrm{max}}$ increases further. As $B_{\textrm{D}}^{\textrm{max}}$ increases, the possibility of allocating destination jamming power increases which initially leads to better energy efficiency due to increased expected total transmitted secure bits, however, if jamming power further increases, energy efficiency decreases due to increased total power expenditure.
(a) Expected total discounted reward vs $\alpha$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(b) Energy efficiency vs $\alpha$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
Figure 5: Expected total discounted reward and energy efficiency vs self-interference.
Fig 5(a) shows the expected total discounted reward versus $\alpha$ for the algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency is plotted in Fig 5(b). Both Figs 5(a) and 5(b) shows the similar trend with $\alpha$. When $\alpha$ is sufficiently small, we do not see any impact of $\alpha$ on the expected total discounted reward or energy efficiency because at very small $\alpha$, self-interference is negligible as compared to the AWGN at node D. However, when $\alpha$ is large enough, the performance degrades as the influence of self-interference on node D becomes comparable to that of the AWGN.
Figure 6: Energy efficiency vs $\Gamma$ for OJPA algorithm and two IPA algorithms IJPA and ITPA for different $p$ and $q$ with $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
@@
Between the IJPA and ITPA algorithms, the IJPA algorithm is more energy efficient than the ITPA algorithm. The ITPA algorithm optimizes transmit power while keeping jamming power fixed. Whereas, the IJPA algorithm optimizes jamming power with a fixed transmit power. As we maximize the expected total transmitted secure bits, the ITPA algorithm leads to more energy consumption due to fixed jamming power since fixing a particular jamming power only decreases the eavesdropping rate not the useful data rate. In contrast, by utilizing a fixed transmit power and allocating optimal energy for jamming in the IJPA algorithm, the useful data transmission improves, leading to improved energy efficiency for the IJPA algorithm. We also find that the energy efficiency decreases with the increasing EH probability for the OJPA and IJPA algorithms. This observation is similar to that in Fig. 3(b) for the OJPA algorithm. In the case of the ITPA algorithm, the observation is just the opposite of the OJPA and IJPA algorithms. Though the IJPA algorithm is more energy efficient than the ITPA algorithm, the ITPA algorithm can take better advantage of increased EH probability than the IJPA algorithm as the energy efficiency of the ITPA algorithm improves with the increasing EH probability.
@@ VIII Conclusion
In this paper, we consider a wireless network with a source and a destination in the presence of an eavesdropper where both the source and the destination are equipped with EH devices with limited battery and the destination has full-duplex jamming capability. We study the problem of joint transmit and jamming power allocation at the source and the destination, respectively, to maximize the long-term expected total transmitted secure bits until the network stops functioning. We formulate infinite-horizon Markov decision process problems for the joint optimization solutions. An optimal algorithm OJPA and a sub-optimal algorithm RSJPA are proposed using the PI algorithm in the RL framework. The results are compared with other sub-optimal algorithms, i.e., GA and NA. Computational complexities of the joint power allocation algorithms are provided. We observe that the proposed RSJPA algorithm achieves nearly optimal secrecy performance with significantly less computational complexity than the OJPA algorithm as it adopts a hybrid approach between the OJPA algorithm and the GA. When we compare the energy efficiency of the OJPA algorithm with two individual power allocation algorithms ITPA and IJPA, and all the sub-optimal joint power allocation algorithms RSJPA, GA, and NA, the OJPA performs the best. Hence, the OJPA algorithm not only provides the best secrecy performance, but also the most energy efficient. Furthermore, the RSJPA algorithm, though sub-optimal, can be a balanced choice between the computational complexity and secrecy performance for joint power allocation. It is observed that the RSJPA algorithm can reduce computational time by around 75 percent as compared to the OJPA algorithm.

### v2

@@ VII Results and Discussions
Fig. 4(a) plots the expected total discounted reward versus the EH probability $p$ while $q=0.5$ and compares the performance of algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency plot can be found in Fig. 4(b). The same plots of Fig. 4(a) and Fig. 4(b) are replicated in Fig. 4(c) and Fig. 4(d), respectively, when the EH probability $q$ is varied while $p=0.5$. From Fig. 4(a) we find that the expected total discounted reward increases for all of the algorithms as EH probability $p$ increases. However, in Fig. 4(c), the expected total discounted reward increases for all of the algorithms except for the GA algorithm as EH probability $q$ increases. A higher probability of EH leads to more energy available at S or D which leads to more expected total transmitted secure bits until the network stops functioning, hence, the observation except for the GA in Fig. 4(c). We also observe that the rate of performance improvement with the EH probability is higher in Fig. 4(a) when $p$ increases at S as compared to Fig. 4(c) when $q$ increases at D.
More expected total transmitted secure bits do not mean more energy efficiency. Energy efficiency decreases with the increasing EH probability $p$ in Fig. 4(b) for the OJPA algorithm and all the algorithms in Fig. 4(d) with the increasing EH probability $q$. However, for the RSJPA, GA, and NA algorithms, the energy efficiency increases in Fig. 4(b). As energy efficiency is the ratio of total transmitted secure bits and total energy expenditure in the system, it seems for the algorithms the rate of increase in the total transmitted secure bits and the rate of increase in the energy expenditure are not the same in various parameter combinations. This suggests that the optimization for total transmitted secure bits alone without taking energy efficiency into account is not beneficial. Rather one needs to optimize the system taking both into account.
@@
(a) Expected total discounted reward versus $\alpha$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
(b) Energy efficiency versus $\alpha$ for joint power allocation algorithms where $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
Figure 6: Expected total discounted reward and energy efficiency versus self-interference.
Figs. 5(a) and 5(c) plot expected total discounted reward versus $B_{\textrm{S}}^{\textrm{max}}$ and $B_{\textrm{D}}^{\textrm{max}}$, respectively and compares the performance of algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency plots are shown in Figs. 5(b) and 5(d), respectively. We observe that as $B_{\textrm{S}}^{\textrm{max}}$ or $B_{\textrm{D}}^{\textrm{max}}$ increases, the expected total discounted reward improves. When $B_{\textrm{S}}^{\textrm{max}}$ and $B_{\textrm{D}}^{\textrm{max}}$ is larger, nodes S and D can store more harvested energy which leads to more expected total transmitted secure bits. If we consider energy efficiency in Figs. 5(b) and 5(d), energy efficiency also improves when $B_{\textrm{S}}^{\textrm{max}}$ increases, however, the same observation is not true when $B_{\textrm{D}}^{\textrm{max}}$ increases. When $B_{\textrm{D}}^{\textrm{max}}$ increases, energy efficiency first improves, however, later degrades as $B_{\textrm{D}}^{\textrm{max}}$ increases further. As $B_{\textrm{D}}^{\textrm{max}}$ increases, the possibility of allocating destination jamming power increases which initially leads to better energy efficiency due to increased expected total transmitted secure bits, however, if jamming power further increases, energy efficiency decreases due to increased total power expenditure.
Fig. 6(a) shows the expected total discounted reward versus $\alpha$ for the algorithms OJPA, RSJPA, GA, and NA. The corresponding energy efficiency is plotted in Fig 6(b). Both Figs 6(a) and 6(b) shows the similar trend with $\alpha$. When $\alpha$ is sufficiently small, we do not see any impact of $\alpha$ on the expected total discounted reward or energy efficiency because at very small $\alpha$, self-interference is negligible as compared to the AWGN at node D. However, when $\alpha$ is large enough, the performance degrades as the influence of self-interference on node D becomes comparable to that of the AWGN.
Figure 7: Expected total discounted reward versus $\Gamma$ when $p=q=0.5$.
[…]
Figure 8: Energy efficiency versus $\Gamma$ for OJPA algorithm and two IPA algorithms IJPA and ITPA for different $p$ and $q$ with $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
@@
Between the IJPA and ITPA algorithms, the IJPA algorithm is more energy efficient than the ITPA algorithm. The ITPA algorithm optimizes transmit power while keeping jamming power fixed. Whereas, the IJPA algorithm optimizes jamming power with a fixed transmit power. As we maximize the expected total transmitted secure bits, the ITPA algorithm leads to more energy consumption due to fixed jamming power since fixing a particular jamming power only decreases the eavesdropping rate not the useful data rate. In contrast, by utilizing a fixed transmit power and allocating optimal energy for jamming in the IJPA algorithm, the useful data transmission improves, leading to improved energy efficiency for the IJPA algorithm. We also find that the energy efficiency decreases with the increasing EH probability for the OJPA and IJPA algorithms. This observation is similar to that in Fig. 4 for the OJPA algorithm. In the case of the ITPA algorithm, the observation is just the opposite of the OJPA and IJPA algorithms. Though the IJPA algorithm is more energy efficient than the ITPA algorithm, the ITPA algorithm can take better advantage of increased EH probability than the IJPA algorithm as the energy efficiency of the ITPA algorithm improves with the increasing EH probability.
Figure 9: Expected total reward versus average harvested energy $\bar{H}_{\textrm{S}}(\bar{H}_{\textrm{D}})$ for OJPA and BI algorithms with $B_{\textrm{S}}^{\textrm{max}}=B_{\textrm{D}}^{\textrm{max}}=5$ energy units.
To check how optimal policy changes for various timelines, we compare the performance of our system, where the network lifetime $K$ is a random variable, with that of the same system where the network lifetime is a finite constant. In the former case, the joint transmit and jamming power optimization is formulated as an infinite-horizon MDP problem, whereas in the latter case, the joint transmit and jamming power optimization is formulated as a finite-horizon MDP problem. In a finite-horizon problem, the optimal policy is obtained using the Backward Induction (BI) algorithm [11, 8]. For a fair comparison, we assume the mean of the network lifetime in the OJPA algorithm is the same as the network lifetime in the BI algorithm. In Fig. 9, we examine the expected total reward for both the algorithms versus average harvested energy $\bar{H}_{\textrm{S}}$ and $\bar{H}_{\textrm{D}}$ individually at S and D, respectively. Average harvested energy at S and D are defined as $\bar{H}_{\textrm{S}}=pE_{\textrm{S}}$ and $\bar{H}_{\textrm{D}}=qE_{\textrm{D}}$. Fig. 9 plots two cases in the same figure: i) $\bar{H}_{\textrm{S}}$ is on the $x$-axis with $p=0.5$ when $\bar{H}_{\textrm{D}}=1.6$ $(q=0.8,E_{\textrm{D}}=2)$ and ii) $\bar{H}_{\textrm{D}}$ is on the $x$-axis with $q=0.5$ when $\bar{H}_{\textrm{S}}=1.6$ $(p=0.8,E_{\textrm{S}}=2)$. We note that in the $y$-axis, we plot the expected total reward instead of the expected total discounted reward as the network lifetime is known. The plots are also shown in two network lifetime scenarios when $K=10$ and $K=20$.
We observe from Fig. 9 that the OJPA algorithm performs better than the BI algorithm. This is because the BI algorithm does not take into account the randomness of the network lifetime which OJPA does. We also notice that the expected total reward saturates after the average harvested energy reaches a certain threshold. The saturation occurs when the harvested energy exceeds the battery capacity of either S or D, and the surplus harvested energy is lost due to battery overflow. Similar to Fig. 3, we observe that the increase in harvested energy at S is more beneficial. We also notice that the expected total reward increases as the network lifetime increases from $K=10$ to $K=20$, however, the nature of the graphs remains the same.
@@ VIII Conclusion
In this paper, we consider a wireless network with a source and a destination in the presence of an eavesdropper where both the source and the destination are equipped with EH devices with limited battery and the destination has full-duplex jamming capability. We study the problem of joint transmit and jamming power allocation at the source and the destination, respectively, to maximize the long-term expected total transmitted secure bits until the network stops functioning. We formulate infinite-horizon Markov decision process problems for the joint optimization solutions. An optimal algorithm OJPA and a sub-optimal algorithm RSJPA are proposed using the PI algorithm in the RL framework. The results are compared with other sub-optimal algorithms, i.e., GA and NA. Computational complexities of the joint power allocation algorithms are provided. We observe that the proposed RSJPA algorithm achieves nearly optimal secrecy performance with significantly less computational complexity than the OJPA algorithm as it adopts a hybrid approach between the OJPA algorithm and the GA. When we compare the energy efficiency of the OJPA algorithm with two individual power allocation algorithms ITPA and IJPA, and all the sub-optimal joint power allocation algorithms RSJPA, GA, and NA, the OJPA performs the best. Hence, the OJPA algorithm not only provides the best secrecy performance, but also the most energy efficient. The secrecy performance of the OJPA algorithm is also compared with the GRLA and BI algorithms, where OJPA achieves the best performance. Furthermore, the RSJPA algorithm, though sub-optimal, can be a balanced choice between the computational complexity and secrecy performance for joint power allocation. It is observed that the RSJPA algorithm with considering only 505050 percent of total number of states can reduce computational time by around 757575 percent as compared to the OJPA algorithm.